An arch can be modelled by the function: h=0.06d^2+2.13.
At a horizontal distance of 2 m from one end of the arch, how high is the arch, to the nearest tenth of a metre?

Answer by nerdybill(6962)

An arch can be modelled by the function: h=0.06d^2+2.13.
At a horizontal distance of 2 m from one end of the arch, how high is the arch, to the nearest tenth of a metre?
set d to 2 and solve for h:
h=0.06d^2+2.13
h=0.06(2)^2+2.13
h=0.06(4)+2.13
h=0.24+2.13
h=2.37
rounded
h=2.4 m